# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@ 3.2.2 Tier 2: Multi-dimensional parameter metric
The second tier consists of an analysis of the posterior distribution of parameter duplicates in the split cosmological model. Given that the posterior distribution for several sampling parameters is prior-dominated, we restrict the tier 2 test to a subset of parameters $\bm{\theta}$ that are constrained by the data while marginalising over the remaining parameters $\bm{\theta}^{\rm marg}$. We calculate the difference
@@
for each data point in the chain, where $\bm{\theta}_{\rm 1/2}$ denotes the two parameter instances in the subspace of parameters of interest. We then analyse the posterior $P(\Delta\bm{\theta})$ in the subspace of parameters of interest. The posterior of parameter differences is given by
@@
where $P^{\rm marg}({\bm{\theta}}_{1},{\bm{\theta}}_{2})$ denotes the posterior distribution marginalised over the remaining (unconstrained) parameters. In the absence of internal tension in the data, we expect $P(\Delta{\bm{\theta}})$ to be centred on the origin, while internal inconsistencies may shift the posterior away from the origin. To quantify the deviation of the posterior from zero, we follow the approach of Köhlinger et al. (2019) and model the posterior with a kernel density estimator (KDE). We evaluate the KDE at the origin and determine the volume of the posterior where the probability of a shift is higher than the probability of no shift. Mathematically, this is equivalent to
$V=\int\limits_{P(\Delta{\bm{\theta}})<P(0)}\!\!\!\!\!\!{\rm d}\Delta{\bm{\theta}}P(\Delta{\bm{\theta}})\;,$ (13)
while in practice we compute the fraction of (weighted) samples with a posterior value lower than $P(0)$. The significance of the shift $m\sigma$ is computed by identifying $V$ with the probability mass of a one-dimensional Gaussian distribution outside of the interval $[-m\sigma,m\sigma]$. Thus, the tension in levels of sigma is given by
@@ 3.2.3 Tier 3: Posterior predictive distribution metric
The third tier of consistency tests compares the observed data with predictions generated from the posterior distribution of model parameters. This is usually probed via the posterior predictive distribution (PPD), which describes the distribution of data realisations given the observed data and assuming a particular model. Given a set of observed data ${\bm{d}}$ and a model $\mathcal{M}$ the distribution of data realisations $\hat{\bm{d}}$ is given by
@@
Considering the split cosmological model, we infer the TPD for each set of cosmological parameters, ${\bm{\theta}}_{\rm A}$ and ${\bm{\theta}}_{\rm B}$ consisting of theory predictions for the full data vector, ${\bm{t}}({\bm{\theta}}_{\rm A})$ and ${\bm{t}}({\bm{\theta}}_{\rm B})$. We then quantify to what extent the observed data in one subset ${\bm{d}}_{\rm A}$ is compatible with the TPD inferred from the other subset and vice versa. To do so, we draw a realisation of ${\bm{d}}_{\rm A}$ conditioned on the TPD of subset B for each sample, denoted by ${\bm{d}}^{\rm sim}_{\rm A}$. Since the conditional distribution of one set of variables conditioned on the other is Gaussian if both sets are jointly Gaussian, the simulated data points are given by a multivariate Gaussian distribution $\mathcal{N}(\mu^{\rm sim}_{\rm A},\mathbf{C}^{\rm sim}_{\rm A})$ with (see for example Bishop, 2006)
$\mu^{\rm sim}_{\rm A}$ $={\bm{t}}_{\rm A}({\bm{\theta}}_{\rm B})+\mathbf{C}_{\rm AB}\mathbf{C}^{-1}_{\rm BB}\left[{\bm{d}}_{\rm B}-{\bm{t}}_{\rm B}({\bm{\theta}}_{\rm B})\right]$ (16)
@@
We quantify the consistency between data regions in terms of the $p$-value $p(A|B)$, which is given by the fraction of posterior samples with
@@
The $p$-value quantifies the probability of the data in data subset A being a realisation of the TPD of subset B. Thus, low $p$-values indicate an internal inconsistency of the data. We note that Doux et al. (2021) show that this method of quantifying consistencies can result in $p$-values that are biased low if the two posteriors prefer vastly dissimilar regions in parameter space. This can be circumvented by calibrating the $p$-value with simulated data. In this work, we interpret the TPD based $p$-value as a conservative metric for the internal consistency since it can exaggerate a potential tension in the data set and reserve a further calibration of the $p$-value for cases for which it fails to pass our adopted threshold for internal consistency.
@@ 4 Results of internal consistency tests
In this section, we present the results of our internal consistency analysis. We adopt the fiducial CosmoPipe pipeline and sample the parameter space via the Nautilus [^4] https://github.com/johannesulf/nautilus/tree/v0.7.2 sampler (Lange, 2023), which is interfaced with the cosmological parameter estimation code Cosmosis (Zuntz et al., 2015). A list of model parameters is provided in Table 1. As is standard practice in stage-III cosmic shear analyses, we adopt a multivariate Gaussian likelihood. In the split cosmological analysis, we duplicate parameters with uniform priors, while parameters with informative Gaussian priors are not duplicated. An exception is the colour-based split of the catalogue, for which we calibrate separate redshift nuisance parameters and allow for different intrinsic galaxy alignments between the subsets. We model the theoretical prediction for the observed signal in each subset with the two independent sets of cosmological parameters. When comparing theory and data, both subsets are linked through the data covariance matrix, which is computed analytically using the OneCovariance [^5] https://github.com/rreischke/OneCovariance/tree/v1.0.0 code (Reischke et al., 2024). We note that our covariance model adopts the NLA model of intrinsic alignments as opposed to the fiducial NLA-$M$ model. As was shown by Reischke et al. (2024), however, the contribution of intrinsic alignments has a negligible impact on the covariance of KiDS-Legacy data. In W25, we consider $p$-values of $p>0.01$, corresponding to a $2.36\sigma$ offset, to be consistent with the null hypothesis for systematics tests. Therefore, we adopt the same threshold for the internal consistency tests conducted in this work.
@@
Before unblinding of the KiDS-Legacy catalogue, we conducted the full consistency analysis on one blind for all three summary statistics. The blinding process, adopted from Kuijken et al. (2015), involves the generation of two additional catalogues with systematic differences in the measured galaxy shapes, which result in up to ±$\pm 2\sigma$ shifts in the inferred $S_{8}$. As the consistency tests are not sensitive to the overall $S_{8}$ value we do not expect the blinding to have an impact on the internal consistency of the KiDS-Legacy data set. With the exception of the split by angular scales with 2PCFs reported in Sect. 4.1.3, we found no differences between consistency tests for the three statistics. We therefore limit our analysis to COSEBIs, except for the split of the data vector by scale, for which we employ all three summary statistics. When evaluating the consistency in parameter space, we focus on $\Omega_{\rm m}$ and $S_{8}$, which are the two parameters that are mostly constrained by cosmic shear data. Prior to the cosmological analysis, we determine the number of constrained parameters $N_{\Theta}$ that is required for the suspiciousness test, as discussed in Sect. 3.2.1. The results of this analysis are presented in Appendix B. In Appendix C, we present a sensitivity analysis with mock realisations of the fiducial data vector, showing that in the absence of internal tension each metric yields a level of consistency that is compatible with typical noise fluctuations in the data.
@@ 4.1 Splits at the data vector level
In the case of data vector-level splits, we employ the fiducial KiDS-Legacy cosmic shear data vector, covariance matrix, and the prior on the shift in the mean of the redshift distribution per tomographic bin of W25 when conducting the likelihood analysis.
@@
Table 2: Consistency metrics for data vector level splits of KiDS-Legacy data.
@@
 [^6] The first column indicates the summary statistic and the second column lists the split applied to the data vector. The third and fourth columns report the tier 1 evidence-based metric of the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 multi-dimensional parameter metric test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 PPD metric test in terms of the $p$-value for data vector predictions for the data subset listed in the second column inferred from the PPD of the other subset (column 8) and vice versa (column 9).
@@ 4.1.1 Redshift bin split
The first split at the data vector level is designed to test the internal consistency between the six tomographic redshift bins. In this way, we probe the data for any errors in our redshift calibration and redshift-dependent modelling effects, such as the impact of baryon feedback or the effect of IA of galaxies, which has a larger relative contribution to the total signal compared to the lensing signal at low redshifts. For each bin, we divide the theory vector into one subset containing the autocorrelation of the specific bin and its cross-correlation with the remaining redshift bins. The second subset consists of all auto-correlations of the remaining redshift bins and their cross-correlation. This split is analogous to the consistency test between KiDS-1000 redshift bins presented in Asgari et al. (2021). A complementary method of testing the consistency between redshift bins is the entire removal of single tomographic bins from the analysis. This test is demonstrated in W25 and is commonly applied in weak lensing studies (see for example Amon et al., 2022; Li et al., 2023b).
@@
4.1.2 Auto-correlation vs cross-correlation
In addition to the redshift bin split, we apply a split of the data vector between auto- and cross-correlation signals of the tomographic bins. This split allows us to probe the data for systematic effects that affect the two types of correlation signals through different processes. In particular, the individual IA contributions to the cosmic shear signal, given in Eq. (2), can be attributed to either the auto- or the cross-correlation signal. The II term is generated through the alignment of physically close galaxies due to tidal forces of the nearby large-scale structure. Therefore, it predominantly affects the autocorrelation of tomographic bins. The GI term on the other hand is induced by the large-scale structure causing both the intrinsic alignment of nearby galaxies and the lensing of distant galaxies, which leads to an anti-correlation between the shapes of galaxy pairs that are separated in redshift. Therefore, this effect manifests itself primarily in the cross-correlation signal between tomographic bins. We divide the theory vector into one subset consisting of the auto-correlation signal of all six tomographic bins and the second subset containing all cross-correlations between bins.
The bottom right panel of Fig. 1 shows the posterior from the consistency test between the autocorrelation group of all tomographic bins in yellow and the cross-correlation group in red. The constraints from the fiducial cosmic shear analysis are visualised with the black dashed line. The consistency metrics, listed in Table 2 signify an agreement between both groups in all tests. This finding is in agreement with the almost indistinguishable posteriors shown in the last panel of Fig. 1. At first glance, it may be surprising that the auto-correlation and cross-correlation parts of the data vector have the same constraining power as each other, and the fiducial data vector. Typically the GI contribution to the cross-correlation signal provides the main information to constrain the IA amplitude, and as such a cosmic shear analysis that excludes the cross-terms decreases the overall constraining power. The fact that we find the same constraining power can be understood by recognising that the NLA-$M$ intrinsic alignment model parameters in our analysis are prior dominated and also shared between the two data splits with the cross-correlation data informing the auto-correlation IA model. If this was not the case then we would expect to see a degradation in constraining power when analysing the auto- and cross-correlations separately.
@@ 4.1.3 Scales
For the 2PCFs, we split the 9 logarithmically spaced angular bins at an angle of $\theta\approx 18.54\arcmin$, with the first/second set consisting of four/five angular bins ensuring comparable signal-to-noise in both sub-samples [^7] Following previous KiDS analyses, we limit the $\xi_{-}$ correlation function to $\theta>4\arcmin$, removing the first $\theta$-bin from the analysis.. For this comparison, we find a tension at $3.69\sigma$ between small and large scales when considering the suspiciousness test. This can be observed in the left panel of Fig. 2, which shows a preference for lower values of $S_{8}$ at large scales and a preference for higher values at small scales, with a $1.81\sigma$ shift in $S_{8}$ between the two. The PPD-based test, however, concludes that the data vectors are consistent. We note that W25 do not include 2PCFs in the fiducial cosmic shear analysis, only providing 2PCF constraints for completeness and comparison with previous works. This is because the 2PCF is particularly sensitive to the effect of baryon feedback which is challenging to model (see for example Asgari et al., 2020). As we have not optimised the angular scales for a 2PCF KiDS-Legacy analysis (see for example Krause et al., 2021) we conclude that the tension reported by the suspicious tier 1 test is likely to be caused by an insufficient treatment of baryonic effects for the fiducial $\theta\in[2\arcmin,300\arcmin]$ angular scale range. As such we do not expect the cosmological constraints from 2PCFs to be as reliable as those inferred with our fiducial COSEBIs statistic. We further note that on large angular scales, the assumption of a Gaussian likelihood does not hold (see for example Sellentin et al., 2018; Louca & Sellentin, 2020; Joachimi et al., 2021; Oehl & Tröster, 2024), which particularly affects the $\xi_{+}$ correlation function, leading to a potential bias in the inferred $S_{8}$.
@@
With our band power spectra we can test the consistency between small and large multipoles in Fourier space (see fig. 1 of W25 for the band power filter functions that are compact in $\ell$). We divide the eight logarithmically spaced bands between $\ell\in\left[100,1500\right]$ at a limit of $\ell\approx 387$, creating two subsets consisting of four bands each. The middle panel of Fig. 2 shows the cosmological constraints from this split analysis with the second to last row of Table 2 reporting the suite of consistency metrics. We find agreement between the low and high $\ell$ band power measurements for all our tests.
@@
It is hard to define a data split for a scale-sensitivity analysis with COSEBIs as each $E_{n}$ mode is sensitive to a range of $\ell$-scales, varying only in the way those scales are combined (see for example fig. 1 of W25). We therefore chose to conduct a data-split analysis of the COSEBI $n=1$ mode, which carries the majority of the constraining power, versus the other $n=2$ to $n=6$ modes. The result is visualised in the right panel of Fig. 2 with the consistency metrics listed in Table 2, demonstrating an agreement between the COSEBIs modes in all tests.

### v2

@@ 3.2.2 Tier 2: Multi-dimensional parameter metric
The second tier consists of an analysis of the posterior distribution of parameter duplicates in the split cosmological model. Given that the posterior distribution for several sampling parameters is prior-dominated, we restricted the tier 2 test to a subset of parameters, $\bm{\theta}$, that are constrained by the data while marginalising over the remaining parameters $\bm{\theta}^{\rm marg}$. We calculated the difference via
@@
for each data point in the chain, where $\bm{\theta}_{\rm 1/2}$ denotes the two parameter instances in the subspace of parameters of interest. We then analysed the posterior $P(\Delta\bm{\theta})$ in the subspace of parameters of interest. The posterior of parameter differences is given by
@@
where $P^{\rm marg}({\bm{\theta}}_{1},{\bm{\theta}}_{2})$ denotes the posterior distribution marginalised over the remaining (unconstrained) parameters. In the absence of internal tension in the data, we expect $P(\Delta{\bm{\theta}})$ to be centred on the origin, while internal inconsistencies may shift the posterior away from the origin. To quantify the deviation of the posterior from zero, we followed the approach of Köhlinger et al. (2019) and modelled the posterior with a kernel density estimator. We evaluated the kernel density estimator at the origin and determined the volume of the posterior where the probability of a shift is higher than the probability of no shift. Mathematically, this is equivalent to
$V=\int\limits_{P(\Delta{\bm{\theta}})<P(0)}\!\!\!\!\!\!{\rm d}\Delta{\bm{\theta}}P(\Delta{\bm{\theta}}).\;$ (13)
We note that in practice, we compute the fraction of (weighted) samples with a posterior value lower than $P(0)$. The significance of the shift $m\sigma$ is computed by identifying $V$ with the probability mass of a one-dimensional Gaussian distribution outside of the interval $m\sigma,m\sigma]$. Thus, the tension in levels of sigma is given by
@@ 3.2.3 Tier 3: Posterior predictive distribution metric
The third tier of consistency tests compares the observed data with predictions generated from the posterior distribution of model parameters. This is usually probed via the posterior predictive distribution (PPD), which describes the distribution of data realisations given the observed data and assuming a particular model. Given a set of observed data, $\bm{d}$}, and a model, $\mathcal{M}$, the distribution of data realisations, $\hat{\bm{d}}$, is given by
@@
Considering the split cosmological model, we infer the TPD for each set of cosmological parameters, ${\bm{\theta}}_{\rm A}$ and ${\bm{\theta}}_{\rm B}$ consisting of theory predictions for the full data vector, ${\bm{t}}({\bm{\theta}}_{\rm A})$ and ${\bm{t}}({\bm{\theta}}_{\rm B})$. We then quantify to what extent the observed data in one subset ${\bm{d}}_{\rm A}$ is compatible with the TPD inferred from the other subset and vice versa. To do so, we draw a realisation of ${\bm{d}}_{\rm A}$ conditioned on the TPD of subset B for each sample, denoted by ${\bm{d}}^{\rm sim}_{\rm A}$. Since the conditional distribution of one set of variables conditioned on the other is Gaussian if both sets are jointly Gaussian, the simulated data points are given by a multivariate Gaussian distribution $\mathcal{N}(\mu^{\rm sim}_{\rm A},\mathbf{C}^{\rm sim}_{\rm A})$ with (see e.g. Bishop, 2006)
$\mu^{\rm sim}_{\rm A}$ $={\bm{t}}_{\rm A}({\bm{\theta}}_{\rm B})+\mathbf{C}_{\rm AB}\mathbf{C}^{-1}_{\rm BB}\left[{\bm{d}}_{\rm B}-{\bm{t}}_{\rm B}({\bm{\theta}}_{\rm B})\right],$ (16)
@@
We quantified the consistency between data regions in terms of the $p$-value, $p(A|B)$, which is given by the fraction of posterior samples with
@@
The $p$-value quantifies the probability of the data in data subset A being a realisation of the TPD of subset B. Thus, low $p$-values indicate an internal inconsistency of the data. We note that Doux et al. (2021) show that this method of quantifying consistencies can result in $p$-values that are biased low if the two posteriors prefer vastly dissimilar regions in parameter space. This can be circumvented by calibrating the $p$-value with simulated data. In this work, we interpret the TPD-based $p$-value as a conservative metric for the internal consistency since it can exaggerate a potential tension in the dataset and reserve a further calibration of the $p$-value for cases for which it fails to pass our adopted threshold for internal consistency.
@@ 4 Results of internal consistency tests
In this section, we present the results of our internal consistency analysis. We adopted the fiducial CosmoPipe pipeline and sample the parameter space via the Nautilus [^4] https://github.com/johannesulf/nautilus/tree/v0.7.2 sampler (Lange, 2023), interfaced with the cosmological parameter estimation code Cosmosis (Zuntz et al., 2015). A list of model parameters is provided in Table 1. As is standard practice in stage-III cosmic shear analyses, we adopted a multivariate Gaussian likelihood. In the split cosmological analysis, we duplicated parameters with uniform priors, while parameters with informative Gaussian priors were not duplicated. An exception is the colour-based split of the catalogue, for which we calibrated separate redshift nuisance parameters and allowed for different intrinsic galaxy alignments between the subsets. We modelled the theoretical prediction for the observed signal in each subset with the two independent sets of cosmological parameters. When comparing theory and data, both subsets were linked through the data covariance matrix, which was computed analytically using the OneCovariance [^5] https://github.com/rreischke/OneCovariance/tree/v1.0.0 code (Reischke et al., 2025). We note that our covariance model adopts the NLA model of intrinsic alignments, as opposed to the fiducial NLA-$M$ model. As was shown by Reischke et al. (2025), however, the contribution of intrinsic alignments has a negligible impact on the covariance of KiDS-Legacy data. In W25, we considered $p$-values of $p>0.01$, corresponding to a $2.36\sigma$ offset, to be consistent with the null hypothesis for systematics tests. Therefore, we adopted the same threshold for the internal consistency tests conducted in this work.
@@
Before the unblinding of the KiDS-Legacy catalogue, we conducted the full consistency analysis on one blind for all three summary statistics. The blinding process, adopted from Kuijken et al. (2015), involves the generation of two additional catalogues with systematic differences in the measured galaxy shapes, which result in up to $\pm 2\sigma$ shifts in the inferred $S_{8}$. As the consistency tests are not sensitive to the overall $S_{8}$ value, we did not expect the blinding to have an impact on the internal consistency of the KiDS-Legacy dataset. With the exception of the split by angular scales with 2PCFs reported in Sect. 4.1.3, we found no differences between consistency tests for the three statistics. We therefore limited our analysis to COSEBIs, except for the split of the data vector by scale, for which we employed all three summary statistics. When evaluating the consistency in parameter space, we focussed on $\Omega_{\rm m}$ and $S_{8}$, which are the two parameters that are mostly constrained by cosmic shear data. Prior to the cosmological analysis, we determined the number of constrained parameters $N_{\Theta}$ that is required for the suspiciousness test, as discussed in Sect. 3.2.1. The results of this analysis are presented in Appendix B. In Appendix C, we present a sensitivity analysis with mock realisations of the fiducial data vector, showing that in the absence of internal tension each metric yields a level of consistency that is compatible with typical noise fluctuations in the data.
@@ 4.1 Splits at the data vector level
In the case of data vector-level splits, we employed the fiducial KiDS-Legacy cosmic shear data vector, covariance matrix, and the prior on the shift in the mean of the redshift distribution per tomographic bin of W25 when conducting the likelihood analysis.
@@
Table 2: Consistency metrics for data vector level splits of KiDS-Legacy data. [^6] The first column indicates the summary statistic and the second column lists the split applied to the data vector. The third and fourth columns report the tier 1 evidence-based metric of the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 multi-dimensional parameter metric test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 PPD metric test in terms of the $p$-value for data vector predictions for the data subset listed in the second column inferred from the PPD of the other subset (column 8) and vice versa (column 9).
@@ 4.1.1 Redshift bin split
The first split at the data vector level is designed to test the internal consistency between the six tomographic redshift bins. In this way, we can probe the data for any errors in our redshift calibration and redshift-dependent modelling effects, such as the impact of baryon feedback or the effect of IA of galaxies, which has a larger relative contribution to the total signal compared to the lensing signal at low redshifts. For each bin, we divided the theory vector into one subset containing the autocorrelation of the specific bin and its cross-correlation with the remaining redshift bins. The second subset consisted of all auto-correlations of the remaining redshift bins and their cross-correlation. This split is analogous to the consistency test between KiDS-1000 redshift bins presented in Asgari et al. (2021). A complementary method of testing the consistency between redshift bins is the entire removal of single tomographic bins from the analysis. This test is demonstrated in W25 and is commonly applied in weak lensing studies (see for example Amon et al., 2022; Li et al., 2023b).
@@
4.1.2 Auto-correlation versus cross-correlation
In addition to the redshift bin split, we applied a split of the data vector between auto- and cross-correlation signals of the tomographic bins. This split allowed us to probe the data for systematic effects that affect the two types of correlation signals through different processes. In particular, the individual IA contributions to the cosmic shear signal, given in Eq. (2), can be attributed to either the auto- or the cross-correlation signal. The II term is generated through the alignment of physically close galaxies due to tidal forces of the nearby large-scale structure. Therefore, it predominantly affects the autocorrelation of tomographic bins. The GI term on the other hand is induced by the large-scale structure causing both the intrinsic alignment of nearby galaxies and the lensing of distant galaxies, which leads to an anti-correlation between the shapes of galaxy pairs that are separated in redshift. Therefore, this effect manifests itself primarily in the cross-correlation signal between tomographic bins. We divided the theory vector into one subset consisting of the auto-correlation signal of all six tomographic bins and the second subset containing all cross-correlations between bins.
The bottom right panel of Fig. 1 shows the posterior from the consistency test between the autocorrelation group of all tomographic bins in yellow and the cross-correlation group in red. The constraints from the fiducial cosmic shear analysis are visualised with the black dashed line. The consistency metrics, listed in Table 2 signify an agreement between both groups in all tests. This finding is in agreement with the almost indistinguishable posteriors shown in the last panel of Fig. 1. At first glance, it may be surprising that the auto-correlation and cross-correlation parts of the data vector have the same constraining power as each other, and the fiducial data vector. Typically the GI contribution to the cross-correlation signal provides the main information to constrain the IA amplitude, and as such a cosmic shear analysis that excludes the cross-terms decreases the overall constraining power. The fact that we have found the same constraining power can be understood by recognising that the NLA-$M$ intrinsic alignment model parameters in our analysis are prior dominated and also shared between the two data splits with the cross-correlation data informing the auto-correlation IA model. If this was not the case then we would expect to see a degradation in constraining power when analysing the auto- and cross-correlations separately.
@@ 4.1.3 Scales
For the 2PCFs, we split the nine logarithmically spaced angular bins at an angle of $\theta\approx 18.54\arcmin$, with the first and second set consisting of four and five angular bins, respectively, ensuring comparable signal-to-noise in both sub-samples [^7] Following previous KiDS analyses, we limit the $\xi_{-}$ correlation function to $\theta>4\arcmin$, removing the first $\theta$-bin from the analysis.. For this comparison, we found tension at $3.69\sigma$ between small and large scales when considering the suspiciousness test. This can be observed in the left panel of Fig. 2, which shows a preference for lower values of $S_{8}$ at large scales and a preference for higher values at small scales, with a $1.81\sigma$ shift in $S_{8}$ between the two. The PPD-based test, however, concludes that the data vectors are consistent. We note that W25 do not include 2PCFs in the fiducial cosmic shear analysis, only providing 2PCF constraints for completeness and comparison with previous works. This is because the 2PCF is particularly sensitive to the effect of baryon feedback which is challenging to model (see for example Asgari et al., 2020). As we have not optimised the angular scales for a 2PCF KiDS-Legacy analysis (see for example Krause et al., 2021) we conclude that the tension reported by the suspicious tier 1 test is likely to be caused by an imperfect modelling of baryonic effects for the fiducial $\theta\in[2\arcmin,300\arcmin]$ angular scale range. As such we do not expect the cosmological constraints from 2PCFs to be as reliable as those inferred with our fiducial COSEBIs statistic, which restrict the range of physical scales entering the analysis (see for example fig. 1 in W25), making them less susceptible to scale-dependent effects. We further note that on large angular scales, the assumption of a Gaussian likelihood does not hold (see for example Sellentin et al., 2018; Louca & Sellentin, 2020; Joachimi et al., 2021; Oehl & Tröster, 2024), which particularly affects the $\xi_{+}$ correlation function, leading to a potential bias in the inferred $S_{8}$.
@@
With our band power spectra we can test the consistency between small and large multipoles in Fourier space (see fig. 1 in W25 for the band power filter functions that are compact in $\ell$). We divide the eight logarithmically spaced bands between $\ell\in\left[100,1500\right]$ at a limit of $\ell\approx 387$, creating two subsets consisting of four bands each. The middle panel of Fig. 2 shows the cosmological constraints from this split analysis with the second to last row of Table 2 reporting the suite of consistency metrics. We found an agreement between the low and high $\ell$ band power measurements for all our tests.
@@
It is hard to define a data split for a scale-sensitivity analysis with COSEBIs as each $E_{n}$ mode is sensitive to a range of $\ell$-scales, varying only in the way those scales are combined (see for example fig. 1 in W25). We therefore chose to conduct a data-split analysis of the COSEBI $n=1$ mode, which carries the majority of the constraining power, versus the other $n=2$ to $n=6$ modes. The result is visualised in the right panel of Fig. 2 with the consistency metrics listed in Table 2, demonstrating an agreement between the COSEBIs modes in all tests.